OpenStudy (anonymous):

One point on the line y=2x+3 has coordinates:

OpenStudy (shamim):

If x=0, y=?

TheSmartOne (thesmartone):

There could be millions of coordinates on the line... Do what @shamim said and you can find the y-intercept of the line :)

OpenStudy (anonymous):

just put a value for x and find the corresponding value of y like if x=1 y=2+3 if x=2 y=4+7 and so on..